Calculate the [H3O+], [OH], pH, and pOH for a solution prepared by dissolving 3.1g KOH(s) (FW=56g/mol) unto enough water to give a volume of 0.500L.
1
Expert's answer
2019-11-21T06:25:47-0500

Answer on Question #99054 – Chemistry – Other


Task:

Calculate the [H3O+], [OH-], pH, and pOH for a solution prepared by dissolving 3.1g KOH(s) (FW=56 g/mol) unto enough water to give a volume of 0.500L.


Solution:

1) Find the molar concentration of the KOH:

Cm = n/V = m / (M*V).

Then,

C(KOH) = m(KOH) / (M(KOH)*V) = 3.1 / (56*0.5) = 3.1 / 28 = 0.1107 (mol/L).


2) dissociation reaction of KOH:

KOH = K+ + OH-;

KOH is strong electrolyte. Then,

C(KOH) = [OH-] = 0.1107 mol/L;


3)pOH = -log[OH-] = 0.96;

pOH = 0.96.


4) pOH + pH = 14;

pH = 14- pOH = 14 - 0.96 = 13.04;

pH = 13.04.


5) pH = -log[H3O+];

[H3O+] = 10-pH = 9.12*10-14 mol/L;

[H3O+] = [H+] = 9.12*10-14 mol/L.


Answer: [H3O+] = 9.12*10-14 mol/L ; [OH-] = 0.1107 mol/L; pH = 13.04; pOH = 0.96.